OpenStudy (anonymous):

What is the equation of a line passing through (-8, -3) and having a slope of -4? y = −4x + 20 y = 4x + 20 y = −4x – 35 y = 4x – 35 y = -4x + 4

OpenStudy (anonymous):

well u can use this formula : y = m(x-x1) + y1 which (x1,y1) is your point and m is slope.

OpenStudy (anonymous):

i have but ive gotten the past 3 questions wrong,i dont think im doing it right

OpenStudy (anonymous):

i'll help you...now x1 = -8 and y1=-3 so : y = -4(x+8) - 3 = -4x - 32 + -3 = -4x - 35

OpenStudy (anonymous):

u can see that 3rd one is the answer ;)

OpenStudy (anonymous):

It was :),thank you for your help.

OpenStudy (anonymous):

you're welcome ;)